what is the equation used to figure out
-If Tanisha has $100 to invest at 7% per annum compounded quarterly, how long will be before she has $150?
-Compounding quarterly
-Compounding continuously

Answer by Boreal(15235)

150=100(1+.07/4)^4n is the formula, the 4n being the number of compoundings in 4 years
1.5=(1+.0175)^4n, dividing by 100
ln 1.5=4n*ln(1.0175)
4n=ln(1.5)/ln(1.0175)=23.37
n is about 6 years (5.84) (on the 24 th compounding)
============
1.5=e^(.07*t)
ln(1.5)=0.07t
t=ln (1.5)/0.07=5.79 years